# From Equations to Insights: Unraveling Symbolic Structures in PDEs with LLMs

### v2

@@ C.3 Detailed computational results for FEX
True solution Iter (Orig) Iter (LLM) Time (Orig) Time (LLM) Err (Orig) Err (LLM) $8\,x_{1}\,\mathrm{exp}(x_{2})$ 42.4 1.2 74.44 0.5900 $3.75\times 10^{-4}$ $4.91\times 10^{-4}$ $2128\,x_{1}^{12}+$ 2 40.2 1.2 70.704 0.516 $3.97\times 10^{-2}$ $2.14\times 10^{-2}$ $-32\,x_{2}^{10}$ 35.2 1.2 71.586 0.536 $7.44\times 10^{-3}$ $6.71\times 10^{-3}$ $-8\,x_{1}^{4}\,x_{2}$ 25.0 1.2 53.342 0.510 $9.18\times 10^{-4}$ $3.86\times 10^{-4}$ $2\,\cos\bigl{(}4\,x_{2}^{6}\bigr{)}$ 44.2 3.0 85.792 4.326 $1.68\times 10^{-2}$ $1.23\times 10^{-2}$ $-8\,x_{2}\,\cos(x_{2})+2$ 32.4 2.4 78.076 4.548 $3.78\times 10^{-4}$ $3.06\times 10^{-4}$ $28\,x_{1}^{2}\,\cos(x_{2})+$ 2 42.8 3.2 85.55 6.154 $2.05\times 10^{-4}$ $1.85\times 10^{-4}$ $-4\,\mathrm{exp}(x_{3})+4\,\sin(x_{1})+2$ 20.2 1.6 43.712 2.150 $1.14\times 10^{-4}$ $2.04\times 10^{-4}$ 2($22\bigl{(}2\,x_{2}^{2}-2\,x_{3}^{4}\bigr{)}^{2}+$ 2 31.8 3.0 69.054 3.400 $6.55\times 10^{-3}$ $2.66\times 10^{-3}$ $-32\,\sin(x_{2})^{4}$ 15.6 1.6 36.210 1.704 $4.34\times 10^{-3}$ $4.66\times 10^{-3}$ $-4\,x_{2}^{3}-4\,x_{3}^{4}+16$ 20.6 2.4 45.110 2.716 $3.88\times 10^{-4}$ $4.68\times 10^{-4}$ 2exp($22\,\mathrm{exp}\bigl{(}-2\,\mathrm{exp}(x_{3})+2\,\cos(x_{3})\bigr{)}+$ 2 22.0 2.8 52.848 4.994 $3.83\times 10^{-4}$ $4.28\times 10^{-4}$ $28\,x_{1}^{2}\,x_{3}+$ 2 11.4 1.6 26.022 1.742 $3.58\times 10^{-4}$ $1.82\times 10^{-4}$ $2\,\sin\bigl{(}2\,\sin(x_{1})-2\,\cos(x_{3})\bigr{)}$ 19.2 3.0 47.168 4.768 $7.78\times 10^{-5}$ $3.13\times 10^{-5}$ $4\bigl{(}2\,\cos(x_{2})-2\,\cos(x_{3})\bigr{)}^{4}$ 10.2 1.6 21.852 1.476 $3.80\times 10^{-3}$ $3.12\times 10^{-3}$ 2($22\bigl{(}2\,x_{1}+2\,\mathrm{exp}(x_{1})\bigr{)}^{2}-$ 2 26.4 4.2 44.664 6.108 $1.07\times 10^{-3}$ $2.94\times 10^{-3}$ $22\,\sin\bigl{(}2\,x_{2}^{3}+2\,x_{3}^{4}\bigr{)}+$ 2 29.0 4.8 73.818 12.376 $6.12\times 10^{-4}$ $3.11\times 10^{-4}$ 2($22\bigl{(}-2\,x_{1}+2\,\mathrm{exp}(x_{3})\bigr{)}^{2}-$ 2 20.4 3.4 35.796 4.762 $1.96\times 10^{-3}$ $2.52\times 10^{-3}$ $22\,\cos\bigl{(}2\,\mathrm{exp}(x_{1})+2\,\sin(x_{2})\bigr{)}+$ 2 69.6 11.6 157.764 27.654 $3.92\times 10^{-5}$ $1.17\times 10^{-4}$ $2\,\sin\bigl{(}2\,\sin(x_{2})-2\,\cos(x_{2})\bigr{)}$ 18.2 3.2 37.792 5.254 $1.01\times 10^{-5}$ $3.35\times 10^{-5}$ $2\bigl{(}-2\,x_{2}^{3}+2\,\sin(x_{3})\bigr{)}^{3}$ 18.2 3.2 42.932 4.500 $1.20\times 10^{-2}$ $5.09\times 10^{-3}$ $2-2\,\mathrm{exp}\bigl{(}4\,x_{2}\,\cos(x_{2})\bigr{)}$ 77.2 15.4 160.186 41.838 $2.89\times 10^{-2}$ $2.89\times 10^{-3}$ $8\,x_{2}^{4}-2\bigl{(}2\,\cos(x_{2})+2\bigr{)}^{2}$ 100.0 20.0 231.106 52.544 $9.44\times 10^{-3}$ $2.13\times 10^{-3}$ $-4\,x_{2}^{4}+4\,x_{2}^{3}+2$ 15.8 3.2 31.420 5.098 $5.81\times 10^{-4}$ $1.03\times 10^{-3}$ $2-2\bigl{(}-2\,\mathrm{exp}(x_{1})+2\,\cos(x_{1})\bigr{)}^{2}$ 38.0 7.8 82.448 15.878 $2.88\times 10^{-3}$ $2.21\times 10^{-3}$ $2\,\sin\bigl{(}2\,x_{3}^{3}-2\,\mathrm{exp}(x_{3})\bigr{)}$ 35.6 8.0 83.972 19.326 $1.51\times 10^{-4}$ $4.52\times 10^{-4}$ $-4\,x_{2}^{4}+4\,x_{3}^{4}+8\,\cos(x_{3})$ 25.6 5.8 62.552 13.770 $1.42\times 10^{-3}$ $2.92\times 10^{-4}$ $2\,\sin\bigl{(}2\,x_{3}^{3}+2\,\cos(x_{3})\bigr{)}$ 13.6 3.4 21.496 5.974 $3.35\times 10^{-4}$ $1.46\times 10^{-4}$ $2\,\mathrm{exp}\bigl{(}2\,x_{2}^{2}-2\,\sin(x_{2})\bigr{)}$ 24.2 6.2 62.858 12.180 $2.04\times 10^{-2}$ $1.71\times 10^{-2}$ $64\,\sin(x_{2})^{2}\,\cos(x_{1})^{2}$ 32.8 9.8 78.910 22.980 $1.09\times 10^{-2}$ $2.66\times 10^{-3}$ $4\,\mathrm{exp}\bigl{(}2\,\sin(x_{2})+2\bigr{)}$ 5.2 1.6 8.154 1.280 $5.37\times 10^{-3}$ $7.00\times 10^{-3}$ $512\,\sin(x_{1})^{4}\,\cos(x_{1})^{4}$ 49.4 15.4 113.850 40.554 $3.42\times 10^{-2}$ $2.53\times 10^{-2}$ $2\,\mathrm{exp}(2\,x_{2})-2\,\mathrm{exp}(2\,x_{1}^{2}-2\,x_{2}^{2})$ 70.8 22.4 148.142 49.740 $4.08\times 10^{-2}$ $7.19\times 10^{-3}$ $2\,\mathrm{exp}\bigl{(}2\,x_{3}^{4}-2\,\mathrm{exp}(x_{3})\bigr{)}$ 17.0 5.4 32.546 10.676 $2.73\times 10^{-3}$ $4.07\times 10^{-3}$ $4\bigl{(}2\,x_{2}^{4}+2\,x_{2}^{3}\bigr{)}^{2}$ 17.6 6.0 28.374 11.458 $1.47\times 10^{-2}$ $1.10\times 10^{-2}$ $2\bigl{(}2\,x_{2}^{4}-2\,x_{2}^{2}\bigr{)}^{3}$ 51.6 17.6 106.732 42.008 $2.45\times 10^{-3}$ $1.08\times 10^{-3}$ $-2\bigl{(}2\,x_{1}^{2}-2\bigr{)}^{4}$ 3.4 1.2 4.070 0.554 $8.30\times 10^{-3}$ $1.25\times 10^{-2}$ $-2\,\sin\bigl{(}2\,\sin(x_{2})-2\bigr{)}$ 2.8 1.0 4.670 0.000 $1.72\times 10^{-4}$ $1.49\times 10^{-4}$ $2\,\mathrm{exp}\bigl{(}2\,\cos(x_{2})+2\bigr{)}$ 5.6 2.0 8.260 2.170 $2.47\times 10^{-3}$ $9.79\times 10^{-3}$ $4\,\cos\bigl{(}4\,x_{2}^{4}\bigr{)}\,\cos\bigl{(}2\,x_{1}^{4}-2\bigr{)}$ 40.4 15.8 91.274 38.016 $6.25\times 10^{-3}$ $5.63\times 10^{-3}$ $2\,\sin\bigl{(}4\,\mathrm{exp}(x_{1})\bigr{)}$ 5.8 2.4 9.644 4.926 $2.67\times 10^{-3}$ $1.26\times 10^{-2}$ $2\bigl{(}2\,x_{3}^{4}-2\bigr{)}^{3}+2\,\cos\bigl{(}2\,x_{2}-2\bigr{)}$ 76.8 31.8 175.690 78.252 $1.99\times 10^{-2}$ $5.17\times 10^{-3}$ $16\,\sin(x_{3})$ 2.4 1.0 2.362 0.000 $1.82\times 10^{-4}$ $8.07\times 10^{-4}$ 2exp($22\,\mathrm{exp}\bigl{(}2\,x_{1}^{4}+2\,\sin(x_{3})\bigr{)}+$ 2 24.6 10.6 57.754 22.852 $7.66\times 10^{-3}$ $8.73\times 10^{-3}$ $22\,\sin\bigl{(}2\,x_{1}^{4}-2\,\sin(x_{3})\bigr{)}-$ 2 55.2 24.0 126.722 62.458 $1.90\times 10^{-4}$ $1.30\times 10^{-4}$ $22\,\sin\bigl{(}2\,x_{3}+2\bigr{)}-$ 2 2.2 1.0 3.464 0.000 $1.86\times 10^{-4}$ $4.64\times 10^{-4}$ $2\,\cos\bigl{(}2\,\sin(x_{2})+2\,\cos(x_{3})\bigr{)}$ 8.0 3.8 15.570 6.608 $1.37\times 10^{-4}$ $9.22\times 10^{-5}$ $2\,\cos\bigl{(}2\,x_{1}^{3}-2\,x_{3}^{2}\bigr{)}$ 20.6 10.0 55.294 23.446 $9.84\times 10^{-4}$ $4.13\times 10^{-4}$ 2exp($22\,\mathrm{exp}\bigl{(}2\,x_{1}^{2}+2\,\sin(x_{3})\bigr{)}+$ 2 17.2 8.4 39.844 19.168 $9.55\times 10^{-3}$ $7.52\times 10^{-3}$ $232\,x_{3}^{6}+$ 2 2.0 1.0 2.084 0.000 $1.10\times 10^{-2}$ $2.93\times 10^{-3}$

### v3

@@ Appendix B Detailed computational results for FEX
True solution Iter (Orig) Iter (LLM) Time (Orig) Time (LLM) Err (Orig) Err (LLM) $8\,x_{1}\,\mathrm{exp}(x_{2})$ 42.4 1.2 74.44 0.5900 $3.75\times 10^{-4}$ $4.91\times 10^{-4}$ $128\,x_{1}^{12}+2$ 40.2 1.2 70.704 0.516 $3.97\times 10^{-2}$ $2.14\times 10^{-2}$ $32\,x_{2}^{10}$ 35.2 1.2 71.586 0.536 $7.44\times 10^{-3}$ $6.71\times 10^{-3}$ $8\,x_{1}^{4}\,x_{2}$ 25.0 1.2 53.342 0.510 $9.18\times 10^{-4}$ $3.86\times 10^{-4}$ $2\,\cos\bigl(4\,x_{2}^{6}\bigr)$ 44.2 3.0 85.792 4.326 $1.68\times 10^{-2}$ $1.23\times 10^{-2}$ $8\,x_{2}\,\cos(x_{2})+2$ 32.4 2.4 78.076 4.548 $3.78\times 10^{-4}$ $3.06\times 10^{-4}$ $8\,x_{1}^{2}\,\cos(x_{2})+2$ 42.8 3.2 85.55 6.154 $2.05\times 10^{-4}$ $1.85\times 10^{-4}$ $4\,\mathrm{exp}(x_{3})+4\,\sin(x_{1})+2$ 20.2 1.6 43.712 2.150 $1.14\times 10^{-4}$ $2.04\times 10^{-4}$ $2\bigl(2\,x_{2}^{2}-2\,x_{3}^{4}\bigr)^{2}+2$ 31.8 3.0 69.054 3.400 $6.55\times 10^{-3}$ $2.66\times 10^{-3}$ $32\,\sin(x_{2})^{4}$ 15.6 1.6 36.210 1.704 $4.34\times 10^{-3}$ $4.66\times 10^{-3}$ $4\,x_{2}^{3}-4\,x_{3}^{4}+16$ 20.6 2.4 45.110 2.716 $3.88\times 10^{-4}$ $4.68\times 10^{-4}$ $\mathrm{exp}\bigl(-2\,\mathrm{exp}(x_{3})+2\,\cos(x_{3})\bigr)+2 22$.0 2.8 52.848 4.994 $3.83\times 10^{-4}$ $4.28\times 10^{-4}$ $8\,x_{1}^{2}\,x_{3}+2$ 11.4 1.6 26.022 1.742 $3.58\times 10^{-4}$ $1.82\times 10^{-4}$ $2\,\sin\bigl(2\,\sin(x_{1})-2\,\cos(x_{3})\bigr)$ 19.2 3.0 47.168 4.768 $7.78\times 10^{-5}$ $3.13\times 10^{-5}$ $4\bigl(2\,\cos(x_{2})-2\,\cos(x_{3})\bigr)^{4}$ 10.2 1.6 21.852 1.476 $3.80\times 10^{-3}$ $3.12\times 10^{-3}$ $2\bigl(2\,x_{1}+2\,\mathrm{exp}(x_{1})\bigr)^{2}-2$ 26.4 4.2 44.664 6.108 $1.07\times 10^{-3}$ $2.94\times 10^{-3}$ $2\,\sin\bigl(2\,x_{2}^{3}+2\,x_{3}^{4}\bigr)+2$ 29.0 4.8 73.818 12.376 $6.12\times 10^{-4}$ $3.11\times 10^{-4}$ $2\bigl(-2\,x_{1}+2\,\mathrm{exp}(x_{3})\bigr)^{2}-2$ 20.4 3.4 35.796 4.762 $1.96\times 10^{-3}$ $2.52\times 10^{-3}$ $2\,\cos\bigl(2\,\mathrm{exp}(x_{1})+2\,\sin(x_{2})\bigr)+2$ 69.6 11.6 157.764 27.654 $3.92\times 10^{-5}$ $1.17\times 10^{-4}$ $2\,\sin\bigl(2\,\sin(x_{2})-2\,\cos(x_{2})\bigr)$ 18.2 3.2 37.792 5.254 $1.01\times 10^{-5}$ $3.35\times 10^{-5}$ $2\bigl(-2\,x_{2}^{3}+2\,\sin(x_{3})\bigr)^{3}$ 18.2 3.2 42.932 4.500 $1.20\times 10^{-2}$ $5.09\times 10^{-3}$ $2-2\,\mathrm{exp}\bigl(4\,x_{2}\,\cos(x_{2})\bigr)$ 77.2 15.4 160.186 41.838 $2.89\times 10^{-2}$ $2.89\times 10^{-3}$ $8\,x_{2}^{4}-2\bigl(2\,\cos(x_{2})+2\bigr)^{2}$ 100.0 20.0 231.106 52.544 $9.44\times 10^{-3}$ $2.13\times 10^{-3}$ $4\,x_{2}^{4}+4\,x_{2}^{3}+2$ 15.8 3.2 31.420 5.098 $5.81\times 10^{-4}$ $1.03\times 10^{-3}$ $2-2\bigl(-2\,\mathrm{exp}(x_{1})+2\,\cos(x_{1})\bigr)^{2}$ 38.0 7.8 82.448 15.878 $88\times 10^{-3} 2.2$$21\times 10^{-3}$ $2\,\sin\bigl(2\,x_{3}^{3}-2\,\mathrm{exp}(x_{3})\bigr)$ 35.6 8.0 83.972 19.326 $1.51\times 10^{-4}$ $4.52\times 10^{-4}$ $4\,x_{2}^{4}+4\,x_{3}^{4}+8\,\cos(x_{3})$ 25.6 5.8 62.552 13.770 $1.42\times 10^{-3}$ $2.92\times 10^{-4}$ $2\,\sin\bigl(2\,x_{3}^{3}+2\,\cos(x_{3})\bigr)$ 13.6 3.4 21.496 5.974 $3.35\times 10^{-4}$ $1.46\times 10^{-4}$ $2\,\mathrm{exp}\bigl(2\,x_{2}^{2}-2\,\sin(x_{2})\bigr)$ 24.2 6.2 62.858 12.180 $2.04\times 10^{-2}$ $1.71\times 10^{-2}$ $64\,\sin(x_{2})^{2}\,\cos(x_{1})^{2}$ 32.8 9.8 78.910 22.980 $1.09\times 10^{-2}$ $2.66\times 10^{-3}$ $4\,\mathrm{exp}\bigl(2\,\sin(x_{2})+2\bigr)$ 5.2 1.6 8.154 1.280 $5.37\times 10^{-3}$ $7.00\times 10^{-3}$ $512\,\sin(x_{1})^{4}\,\cos(x_{1})^{4}$ 49.4 15.4 113.850 40.554 $3.42\times 10^{-2}$ $2.53\times 10^{-2}$ $2\,\mathrm{exp}(2\,x_{2})-2\,\mathrm{exp}(2\,x_{1}^{2}-2\,x_{2}^{2})$ 70.8 22.4 148.142 49.740 $4.08\times 10^{-2}$ $7.19\times 10^{-3}$ $2\,\mathrm{exp}\bigl(2\,x_{3}^{4}-2\,\mathrm{exp}(x_{3})\bigr)$ 17.0 5.4 32.546 10.676 $2.73\times 10^{-3}$ $4.07\times 10^{-3}$ $4\bigl(2\,x_{2}^{4}+2\,x_{2}^{3}\bigr)^{2}$ 17.6 6.0 28.374 11.458 $47\times 10^{-2} 1.1$$10\times 10^{-2}$ $2\bigl(2\,x_{2}^{4}-2\,x_{2}^{2}\bigr)^{3}$ 51.6 17.6 106.732 42.008 $2.45\times 10^{-3}$ $1.08\times 10^{-3}$ $2\bigl(2\,x_{1}^{2}-2\bigr)^{4}$ 3.4 1.2 4.070 0.554 $8.30\times 10^{-3}$ $1.25\times 10^{-2}$ $2\,\sin\bigl(2\,\sin(x_{2})-2\bigr)$ 2.8 1.0 4.670 0.000 $1.72\times 10^{-4}$ $1.49\times 10^{-4}$ $2\,\mathrm{exp}\bigl(2\,\cos(x_{2})+2\bigr)$ 5.6 2.0 8.260 2.170 $2.47\times 10^{-3}$ $9.79\times 10^{-3}$ $4\,\cos\bigl(4\,x_{2}^{4}\bigr)\,\cos\bigl(2\,x_{1}^{4}-2\bigr)$ 40.4 15.8 91.274 38.016 $6.25\times 10^{-3}$ $5.63\times 10^{-3}$ $2\,\sin\bigl(4\,\mathrm{exp}(x_{1})\bigr)$ 5.8 2.4 9.644 4.926 $2.67\times 10^{-3}$ $1.26\times 10^{-2}$ $2\bigl(2\,x_{3}^{4}-2\bigr)^{3}+2\,\cos\bigl(2\,x_{2}-2\bigr)$ 76.8 31.8 175.690 78.252 $1.99\times 10^{-2}$ $5.17\times 10^{-3}$ $16\,\sin(x_{3})$ 2.4 1.0 2.362 0.000 $1.82\times 10^{-4}$ $8.07\times 10^{-4}$ $2\,\mathrm{exp}\bigl(2\,x_{1}^{4}+2\,\sin(x_{3})\bigr)+2$ 24.6 10.6 57.754 22.852 $7.66\times 10^{-3}$ $8.73\times 10^{-3}$ $2\,\sin\bigl(2\,x_{1}^{4}-2\,\sin(x_{3})\bigr)-2$ 55.2 24.0 126.722 62.458 $1.90\times 10^{-4}$ $1.30\times 10^{-4}$ $\sin\bigl(2\,x_{3}+2\bigr)-2 2.2$ 1.0 3.464 0.000 $1.86\times 10^{-4}$ $4.64\times 10^{-4}$ $2\,\cos\bigl(2\,\sin(x_{2})+2\,\cos(x_{3})\bigr)$ 8.0 3.8 15.570 6.608 $1.37\times 10^{-4}$ $9.22\times 10^{-5}$ $2\,\cos\bigl(2\,x_{1}^{3}-2\,x_{3}^{2}\bigr)$ 20.6 10.0 55.294 23.446 $9.84\times 10^{-4}$ $4.13\times 10^{-4}$ $2\,\mathrm{exp}\bigl(2\,x_{1}^{2}+2\,\sin(x_{3})\bigr)+2$ 17.2 8.4 39.844 19.168 $9.55\times 10^{-3}$ $7.52\times 10^{-3}$ $32\,x_{3}^{6}+2$ 2.0 1.0 2.084 0.000 $1.10\times 10^{-2}$ $2.93\times 10^{-3}$